# Estimating the Level of Dialectness Predicts Interannotator Agreement in Multi-dialect Arabic Datasets

### v2

@@ Abstract
On annotating multi-dialect Arabic datasets, it is common to randomly assign the samples across a pool of native Arabic speakers. Recent analyses recommended routing dialectal samples to native speakers of their respective dialects to build higher-quality datasets. However, automatically identifying the dialect of samples is hard. Moreover, the pool of annotators who are native speakers of specific Arabic dialects might be scarce.
Arabic Level of Dialectness (ALDi) was recently introduced as a quantitative variable that measures how sentences diverge from Standard Arabic. On randomly assigning samples to annotators, we hypothesize that samples of higher ALDi scores are harder to label especially if they are written in dialects that the annotators do not speak. We test this by analyzing the relation between ALDi scores and the annotators’ agreement, on 15 public datasets having raw individual sample annotations for various sentence-classification tasks. We find strong evidence supporting our hypothesis for 11 of them. Consequently, we recommend prioritizing routing samples of high ALDi scores to native speakers of each sample’s dialect, for which the dialect could be automatically identified at higher accuracies.
@@ 1 Introduction
Arabic is spoken natively by over 420 million people and is an official language of 24 countries Bergman and Diab (2022), making it an important language for NLP systems. However, NLP for Arabic faces a major challenge in that user-generated text is typically a mixture of Modern Standard Arabic (MSA)—the standardized variant that is taught in schools and used in official communications and newspapers—and regional variants of Dialectal Arabic (DA), which are used in everyday communications, including both speech and social media text Habash (2010). [^1] Refer to §B of the Appendix for futher discussion about the relationship between MSA and DA. While MSA can be largely understood by most Arabic speakers, the different variants of DA are not always fully mutually intelligible.
Despite this mutual unintelligibility, a common practice when developing datasets for multi-dialect Arabic NLP is to randomly recruit annotators without regard to their dialect.
However, routing dialectal content to speakers of a different dialect for annotation or moderation can present real problems. For example, it has been shown to contribute to unjust online content moderation of DA Business for Social Responsibility (2022), and racially biased toxicity annotation in American English varieties Sap et al. (2022). Two recent studies of multi-dialect DA annotation showed that for annotating hate speech or sarcasm, respectively, annotators were more lenient (for hate speech) and more accurate (for sarcasm) when annotating sentences in their native dialect Bergman and Diab (2022); Abu Farha and Magdy (2022).
The authors of both studies made the same recommendation for creating new Arabic datasets, namely to first identify the dialect of each sample and then route it to appropriate annotators.
@@
We test our hypothesis by investigating the impact of ALDi score on inter-annotator agreement for 15 publicly released datasets annotated for 6 different sentence-classification tasks. [^2] Instructions to replicate the experiments can be accessed through https://github.com/AMR-KELEG/ALDi-and-IAA We confirm that for most tasks and datasets, higher ALDi scores correlate with lower annotator agreement. A notable exception is the dialect identification (DI) task, where higher ALDi scores correlate with higher agreement, presumably because it is easier to identify a single dialect for sentences that are strongly dialectal. This finding is encouraging for annotation routing, since automatic DI systems may also have higher accuracy on these sentences. We conclude that a combination of automatic ALDi scoring, followed by DI and annotator routing only for high-ALDi sentences, is a promising strategy for annotating multi-dialect Arabic datasets.
@@ 2 Methodology
We study the impact of ALDi scores on the annotators’ agreement for publicly released Arabic datasets.
We analyze datasets satisfying the following criteria:
@@
For each dataset, we compute the Arabic Level of Dialectness (ALDi) score for each annotated sample (sentence) using the Sentence-ALDi model Keleg et al. (2023), which returns a score from 0 (MSA/non-dialectal) to 1 (strongly dialectal). To investigate the effect of ALDi on annotator agreement, we bin the samples by their ALDi score into 10 bins of width 0.1. We compute % full agree, the percentage of samples in that bin for which all the annotators agreed on a single label. We employ Pearson’s correlation coefficient to analyze the relation between ALDi (represented by each bin’s midpoint ALDi score) and % full agree, and also report the slope of the best-fitting line as a measure of the effect size [^4] The exact values of the slopes and correlation coefficients depend on the number of bins. However, we got similar qualitative results on using 4 or 20 equal-width bins. 10 bins are enough to check if trends are non-linear while keeping a reasonable number of samples in the smallest bins. We also fitted logistic regression (logreg) models using ALDi as a continuous variable and a binary outcome Full Agreement (Yes/No) for each sample. Both analysis tools reveal similar patterns (see Appendix §C) but the binning method provides useful additional visualization.. As aforementioned, our initial hypothesis is that % full agree negatively correlates with high ALDi scores.
@@ 3 Results and Discussion
Figure 1: Scatter plots showing the relationship between binned ALDi scores (x-axis) and the percentage of samples with full annotator agreement (y-axis). The histogram represents the # of samples per bin (with min and max values for any bin labeled on the right-hand axis). The slope of the best-fitting line ($m$) is shown, and to enable visual comparison of slopes, all plots have the same y-axis scale (possibly shifted up or down), except Mawqif (Sentiment Analysis). Note: Statistically significant (p<0.05) correlation coefficients ($\rho$) are marked with *.
@@
Inspecting the trends depicted in Figure 1, strong negative Pearson’s correlation coefficients exist for 8 out of the 12 datasets for the non-DI tasks (sentiment analysis, sarcasm, hate speech, and stance detection).
Both the trends (quantified by the slope m) and the correlation coefficients for most of the tasks indicate that the percentage of samples for which all the annotators assign the same label decreases as the ALDi scores increase, often by a large margin [^5] Refer to Appendix §E for a possible explanation of the unexpected trends of the ArSAS dataset.. We notice different trends for DI that we elaborate on below.
@@
For the ArSarcasm-v1 and iSarcasm datasets, the set of labels for the DI task includes MSA (i.e.: some sentences in these datasets are not dialectal). For both datasets, one notices high percentages of agreement scores for the bin having ALDi scores $\in[0,0.1]$ (generally agreeing that the label is MSA). The percentages decrease for the few succeeding bins, before rising again for the bins with high ALDi scores. Sentences of high ALDi scores (e.g.: $\in[0.8,1]$) are expected to have multiple dialectal cues, which increases the chance of attributing them to a single dialect. For sentences of intermediate ALDi scores, annotators can agree a sentence is not in MSA, however, they would struggle to determine the dialect of the sentence, which is manifested as having lower percentages of the full agreement for these bins.
@@ 4 Analysis of Trends by Class Label
The declining trend for agreement as ALDi scores increase is consistently salient for the negative class of the different tasks (i.e.: Not Sarcastic, Neutral (sentiment), Non-offensive). One explanation is that agreeing on one of these labels requires fully understanding the sample, which is expected to be harder for non-native speakers of the dialects of sentences with high ALDi scores. Moreover, the general usage of MSA for formal communications, and DA for more personal ones might be biasing the annotators. Further controlled analysis is required to investigate these intuitive explanations.
Conversely, the presence of specific words might be a strong cue for the label of the overall sentence, which might be useful for the positive classes such as (i.e.: Sarcastic, Positive/Negative, Obscene/Offensive). Consequently, annotators might be able to agree on a label for a sentence even if it contains some unintelligible segments. Table B2 shows five examples, in which the annotators can use cues to label the whole sentence. Consider the first offensive sentence in Table LABEL:tab:qualitiative_seducated_guess_examples. The presence of the MSA insult (you insect) is enough to guess that the sentence is offensive, even if the remaining segment is not fully intelligible.
It is hard to know whether the annotators relied on these cues to make their decision or not. We think that the usage of cues is an intuitive explanation that needs further validation. We would also like to Lastly, note that agreeing on a label does not imply it is accurate, especially when in the case of relying on cues for annotation.
@@ 5 Conclusion and Recommendation
Previous research recommended routing samples to native speakers of the different Arabic dialects for higher annotation quality.
Our analysis indicates that a large proportion of 6 datasets are samples with ALDi scores < 0.1, which are expected to be MSA samples that can be routed to speakers of any Arabic dialect. Moreover, the lower agreement scores for samples with high ALDi scores show that extra care should be given to these samples. Dataset creators should prioritize routing them to native speakers of the dialects of these samples, which can be automatically identified at higher accuracy as these samples have more dialectal cues.
@@
Another potential extension to this work is to analyze the inter-annotator disagreement on annotating dialectal data for token-level tasks. To the best of our knowledge, all the publicly available token-level Arabic datasets are built by carefully selecting samples written in specific dialects and recruiting native speakers of each of these dialects to perform the annotation, after closely training them. However, even if a multi-dialect token-level dataset is annotated by randomly assigning the samples to speakers of different dialects, the analysis would require a new model to estimate the level of dialectness on the token level, since the Sentence-ALDi model used here works at the sentence level.
@@ Acknowledgments
This work could not have been done without the help of the datasets’ creators who have kindly agreed to share the individual labels for their datasets’ samples. Thanks Ibrahim Abu Farha, Nora Alturayeif, and Manal Alshehri. We also thank Hamdy Mubarak, Nuha Albadi, Nedjma Ousidhoum, and Hala Mulki for trying to help with finding the individual annotator labels for some of their datasets. Lastly, we appreciate the efforts of the anonymous ARR reviewers, action editors, and area chairs. Thanks for the insightful discussions and valuable suggestions.
@@ Appendix A Detailed Description of the Datasets
Table A1: A detailed description of the distribution of the majority-vote labels and the data/paper discrepancies in the datasets with individual-annotator labels included in our study.
@@ Appendix B Code-mixing between MSA and DA
# ??? # hahaha 0.94 Positive(2x) Negative(1x)
YTCB Hatespeech Detection you scabies 0.81 HateSpeech(3x)
Gulf’s colt 0.98 HateSpeech(3x)
Table B2: Four qualitative samples of high ALDi scores. The underlined segments represent the cues that the annotators might have used to choose a label even if they do not fully understand the sentence. Despite the presence of these cues, the annotators still disagreed on labeling the first two samples. We only provide translations for the underlined cue segments.
@@ Appendix C Discussion about the Analysis
Automated metrics such as Fleiss’ Kappa Fleiss (1971) attempt to measure the inter-annotator agreement (IAA) while accounting for the random agreement/disagreement between annotators. In principle, it might be possible to perform a version of our analysis using Fleiss’ Kappa rather than % full agree as the dependent variable. However, computing Fleiss’ Kappa would require knowledge of the individual annotations for each sample. Such annotations are not available for the ArSAS (Sentiment/Speech Act), DART, and DCD datasets as described in Table A1. Since we wanted to include as many data sets as possible, we used % full agree instead.
@@
Both analysis tools achieve similar results. The same 8 non-DI datasets—ASAD, ArSarcasm-v1 (Sentiment/Sarcasm), Mawqif (Sentiment/Sarcasm/Stance), iSarcasm, and YTCB—have significantly strong negative correlation coefficients as in Figure 1, and statistically significant coefficients for the ALDi variable which are less than -0.2. However,
binning the data allows for visualizing the % Full Agreement as a scatter plot, which can reveal whether the relation between ALDi and the agreement is linear or not, in addition to having a visual way for determining how well the best-fitting line models the data.
@@
MSA samples are over-represented in some of the considered datasets. However, this is generally unproblematic for the analysis, so we opted not to discard MSA samples. For the method described in Section 2, the samples of each bin are independently used to estimate the MLE of full agreement between annotators. For Logistic Regression, the MSA samples whose ALDi scores are nearly equal to 0 are expected to only affect the intercept, without greatly changing the coefficient of the continuous ALDi variable.
@@
Figure C1: For each dataset, plots show the estimated probability of Full-agreement according to each dataset’s fitted logistic regression model. Under each plot, the coefficient of ALDi with its 95% confidence interval is visualized. Nearly all datasets (marked with *) have confidence intervals that do not include zero, meaning the effect of ALDi is statistically significant at $p<0.05$. Negative coefficients indicate that higher ALDi scores predict lower agreement.
@@ Appendix D Trends by Class Label
As mentioned in §4, Figures D2, D3, D4, D5, and D6 visualize the impact of ALDi on the annotator agreement after splitting the samples according to their majority-vote labels. We acknowledge that the number of samples in the bins for some classes is not enough to draw concrete conclusions (e.g.: samples with high ALDi scores for the Neutral class of the ArSAS, and Mawqif datasets as per Figure D3).
@@ Appendix E The Rising Trend of ArSAS
The ArSAS dataset stands out as a dataset with a rising trend for the Speech Act Detection task and a falling trend for the Sentiment Analysis task. Samples of ArSAS were jointly annotated for their sentiment and speech act. Despite having 6 different speech acts, which would arguably make speech act detection harder than sentiment analysis, the Assertion and Expression classes represent 95% of the samples. Looking at their respective trends shown in Figure D5, the two acts show two different behaviors. Most of the assertive samples have ALDi scores <0.2 (arguably all are MSA ones). Moreover, the number of Assestion samples with high ALDi scores is not enough to estimate the % full agree for their respective bins. Conversely, the Expression act shows higher agreement as the ALDi score increases.
The creators of ArSAS noticed that most of the Assertion samples were annotated as Neutral, while most of the Expression samples have polarized sentiment (mostly Negative). The annotators might have treated the Assertion class as the act for Objective sentences, while Expression as the act for Subjective sentences. This is arguably easier than sentiment analysis which might explain why annotators agree more on the Speech Act label than the Sentiment label for the ArSAS dataset. Further analysis is required to explain the trends of this dataset.

### v3

@@ Abstract
On annotating multi-dialect Arabic datasets, it is common to randomly assign the samples across a pool of native Arabic speakers. Recent analyses recommended routing dialectal samples to native speakers of their respective dialects to build higher-quality datasets. However, automatically identifying the dialect of samples is hard. Moreover, the pool of annotators who are native speakers of specific Arabic dialects might be scarce. Arabic Level of Dialectness (ALDi) was recently introduced as a quantitative variable that measures how sentences diverge from Standard Arabic. On randomly assigning samples to annotators, we hypothesize that samples of higher ALDi scores are harder to label especially if they are written in dialects that the annotators do not speak. We test this by analyzing the relation between ALDi scores and the annotators’ agreement, on 15 public datasets having raw individual sample annotations for various sentence-classification tasks. We find strong evidence supporting our hypothesis for 11 of them. Consequently, we recommend prioritizing routing samples of high ALDi scores to native speakers of each sample’s dialect, for which the dialect could be automatically identified at higher accuracies.
@@ 1 Introduction
Arabic is spoken natively by over 420 million people and is an official language of 24 countries Bergman and Diab (2022), making it an important language for NLP systems. However, NLP for Arabic faces a major challenge in that user-generated text is typically a mixture of Modern Standard Arabic (MSA)—the standardized variant that is taught in schools and used in official communications and newspapers—and regional variants of Dialectal Arabic (DA), which are used in everyday communications, including both speech and social media text Habash (2010). [^1] Refer to §B of the Appendix for a further discussion about the relationship between MSA and DA. While MSA can be largely understood by most Arabic speakers, the different variants of DA are not always fully mutually intelligible.
Despite this mutual unintelligibility, a common practice when developing datasets for multi-dialect Arabic NLP is to randomly recruit annotators without regard to their dialect. However, routing dialectal content to speakers of a different dialect for annotation or moderation can present real problems. For example, it has been shown to contribute to unjust online content moderation of DA Business for Social Responsibility (2022), and racially biased toxicity annotation in American English varieties Sap et al. (2022). Two recent studies of multi-dialect DA annotation showed that for annotating hate speech or sarcasm, respectively, annotators were more lenient (for hate speech) and more accurate (for sarcasm) when annotating sentences in their native dialect Bergman and Diab (2022); Abu Farha and Magdy (2022). The authors of both studies made the same recommendation for creating new Arabic datasets, namely to first identify the dialect of each sample and then route it to appropriate annotators.
@@
We test our hypothesis by investigating the impact of ALDi score on interannotator agreement for 15 publicly released datasets annotated for 6 different sentence-classification tasks. [^2] Instructions to replicate the experiments can be accessed through https://github.com/AMR-KELEG/ALDi-and-IAA We confirm that for most tasks and datasets, higher ALDi scores correlate with lower annotator agreement. A notable exception is the dialect identification (DI) task, where higher ALDi scores correlate with higher agreement, presumably because it is easier to identify a single dialect for sentences that are strongly dialectal. This finding is encouraging for annotation routing, since automatic DI systems may also have higher accuracy on these sentences. We conclude that a combination of automatic ALDi scoring, followed by DI and annotator routing only for high-ALDi sentences, is a promising strategy for annotating multi-dialect Arabic datasets.
@@ 2 Methodology
We study the impact of ALDi scores on the annotators’ agreement for publicly released Arabic datasets. We analyze datasets satisfying the following criteria:
@@
For each dataset, we compute the Arabic Level of Dialectness (ALDi) score for each annotated sample (sentence) using the Sentence-ALDi model Keleg et al. (2023), which returns a score from 0 (MSA/non-dialectal) to 1 (strongly dialectal). To investigate the effect of ALDi on annotator agreement, we bin the samples by their ALDi score into 10 bins of width 0.1. We compute % full agree, the percentage of samples in that bin for which all the annotators agreed on a single label. We employ Pearson’s correlation coefficient to analyze the relation between ALDi (represented by each bin’s midpoint ALDi score) and % full agree, and also report the slope of the best-fitting line as a measure of the effect size. [^4] The exact values of the slopes and correlation coefficients depend on the number of bins. However, we got similar qualitative results on using 4 or 20 equal-width bins. 10 bins are enough to check if trends are non-linear while keeping a reasonable number of samples in the smallest bins. We also fitted logistic regression (logreg) models using ALDi as a continuous variable and a binary outcome Full Agreement (Yes/No) for each sample. Both analysis tools reveal similar patterns (See Appendix §C) but the binning method provides useful additional visualization. As aforementioned, our initial hypothesis is that % full agree negatively correlates with high ALDi scores.
@@ 3 Results and Discussion
Figure 1: Scatter plots showing the relationship between binned ALDi scores (x-axis) and the percentage of samples with full annotator agreement (y-axis). The histogram represents the # of samples per bin (with min and max values for any bin labeled on the right-hand axis). The slope of the best-fitting line ($m$) is shown, and to enable visual comparison of slopes, all plots have the same y-axis scale (possibly shifted up or down).
Note: Statistically significant (p<0.05) correlation coefficients ($\rho$) are marked with *.
@@
Inspecting the trends depicted in Figure 1, strong negative Pearson’s correlation coefficients exist for 8 out of the 12 datasets for the non-DI tasks (sentiment analysis, sarcasm, hate speech, and stance detection). Both the trends (quantified by the slope m) and the correlation coefficients for most of the tasks indicate that the percentage of samples for which all the annotators assign the same label decreases as the ALDi scores increase, often by a large margin. [^5] Refer to Appendix §E for a possible explanation of the unexpected trends of the ArSAS dataset. We notice different trends for DI that we elaborate on below.
@@
For the ArSarcasm-v1 and iSarcasm datasets, the set of labels for the DI task includes MSA (i.e., some sentences in these datasets are not dialectal). For both datasets, one notices high percentages of agreement scores for the bin having ALDi scores $\in[0,0.1]$ (generally agreeing that the label is MSA). The percentages decrease for the few succeeding bins, before rising again for the bins with high ALDi scores. Sentences of high ALDi scores (e.g., $\in[0.8,1]$) are expected to have multiple dialectal cues, which increases the chance of attributing them to a single dialect. For sentences of intermediate ALDi scores, annotators can agree a sentence is not in MSA, however, they would struggle to determine the dialect of the sentence, which is manifested as having lower percentages of the full agreement for these bins.
@@ 4 Analysis of Trends by Class Label
The declining trend for agreement as ALDi scores increase is consistently salient for the negative class of the different tasks (i.e., Not Sarcastic, Neutral (sentiment), Non-offensive). One explanation is that agreeing on one of these labels requires fully understanding the sample, which is expected to be harder for non-native speakers of the dialects of sentences with high ALDi scores. Moreover, the general usage of MSA for formal communications, and DA for more personal ones might be biasing the annotators. Further controlled analysis is required to investigate these intuitive explanations.
Conversely, the presence of specific words might be a strong cue for the label of the overall sentence, which might be useful for the positive classes (i.e., Sarcastic, Positive/Negative, Obscene/Offensive). Consequently, annotators might be able to agree on a label for a sentence even if it contains some unintelligible segments. Table 2 shows five examples, in which the annotators can use cues to label the whole sentence. Consider the first offensive sentence in the Table. The presence of the MSA insult (you insect) is enough to guess that the sentence is offensive, even if the remaining segment is not fully intelligible. Lastly, note that agreeing on a label does not imply it is accurate, especially when relying on cues for annotation.
@@ 5 Conclusion and Recommendation
Previous research recommended routing samples to native speakers of the different Arabic dialects for higher annotation quality. Our analysis indicates that a large proportion of 6 datasets are samples with ALDi scores < 0.1, which are expected to be MSA samples that can be routed to speakers of any Arabic dialect. Moreover, the lower agreement scores for samples with high ALDi scores show that extra care should be given to these samples. Dataset creators should prioritize routing high-ALDi samples to native speakers of the dialects of these samples, for which the dialects can be automatically identified at higher accuracy as these samples have more dialectal cues.
@@
Another potential extension to this work is to analyze the interannotator disagreement on annotating dialectal data for token-level tasks. To the best of our knowledge, all the publicly available token-level Arabic datasets are built by carefully selecting samples written in specific dialects and recruiting native speakers of each of these dialects to perform the annotation, after closely training them. However, even if a multi-dialect token-level dataset is annotated by randomly assigning the samples to speakers of different dialects, the analysis would require a new model to estimate the level of dialectness on the token level, since the Sentence-ALDi model used here works at the sentence level.
@@ Acknowledgments
This work could not have been done without the help of the datasets’ creators who have kindly agreed to share the individual labels for their datasets’ samples. Thanks, Ibrahim Abu Farha, Nora Alturayeif, and Manal Alshehri. We also thank Hamdy Mubarak, Nuha Albadi, Nedjma Ousidhoum, and Hala Mulki for trying to help with finding the individual annotator labels for some of their datasets. Lastly, we appreciate the efforts of the anonymous ARR reviewers, action editors, and area chairs. Thanks for the insightful discussions and valuable suggestions.
@@ Appendix A Detailed Description of the Datasets
Table A1: A detailed description of the distribution of the majority-vote labels and the data/paper discrepancies in the datasets with individual annotator labels included in our study.
@@ Appendix C Discussion about the Analysis
Automated metrics such as Fleiss’ Kappa Fleiss (1971) attempt to measure the Interannotator Agreement (IAA) while accounting for the random agreement/disagreement between annotators. In principle, it might be possible to perform a version of our analysis using Fleiss’ Kappa rather than % full agree as the dependent variable. However, computing Fleiss’ Kappa would require knowledge of the individual annotations for each sample. Such annotations are not available for the ArSAS (Sentiment/Speech Act), DART, and DCD datasets as described in Table A1. Since we wanted to include as many datasets as possible, we used % full agree instead.
@@
Both analysis tools (correlation analysis and logistic regression) achieve similar results. The same 8 non-DI datasets—ASAD, ArSarcasm-v1 (Sentiment/Sarcasm), Mawqif (Sentiment/Sarcasm/Stance), iSarcasm, and YTCB—have significantly strong negative correlation coefficients as in Figure 1, and statistically significant coefficients for the ALDi variable which are less than -0.2. However, binning the data allows for visualizing the % full agreement as a scatter plot, which can reveal whether the relation between ALDi and the agreement is linear or not, in addition to having a visual way for determining how well the best-fitting line models the data.
@@
MSA samples are over-represented in some of the considered datasets. However, this is generally unproblematic for the analysis, so we opted not to discard the MSA samples. For the method described in Section 2, the samples of each bin are independently used to estimate the MLE of full agreement between annotators. Therefore, the over-representation of MSA samples in some datasets does not impact our analysis.
@@
Figure C1: For each dataset, plots show the estimated probability of full agreement according to each dataset’s fitted logistic regression model. Under each plot, the coefficient of ALDi with its 95% confidence interval is visualized. Nearly all datasets (marked with *) have confidence intervals that do not include zero, meaning the effect of ALDi is statistically significant at $p<0.05$. Negative coefficients indicate that higher ALDi scores predict lower agreement.
@@ Appendix D Trends by Class Label
As mentioned in §4, Figures D2D3D4D5, and D6 visualize the impact of ALDi on the annotator agreement after splitting the samples according to their majority-vote labels. We acknowledge that the number of samples in the bins for some classes is not enough to draw concrete conclusions (e.g., samples with high ALDi scores for the Neutral class of the ArSAS, and Mawqif datasets as per Figure D3).
@@ Appendix E The Rising Trend of ArSAS
The ArSAS dataset stands out as a dataset with a rising trend for the Speech Act Detection task and a falling trend for the Sentiment Analysis task. Samples of ArSAS were jointly annotated for their sentiment and speech act. Despite having 6 different speech acts, which would arguably make speech act detection harder than sentiment analysis, the Assertion and Expression classes represent 95% of the samples. Looking at their respective trends shown in Figure D5, the two acts show two different behaviors. Most of the assertive samples have ALDi scores <0.2 (arguably, all are MSA ones). Moreover, the number of Assestion samples with high ALDi scores is not enough to estimate the % full agree for their respective bins. Conversely, the Expression act shows higher agreement as the ALDi score increases.
The creators of ArSAS noticed that most of the Assertion samples were annotated as Neutral, while most of the Expression samples had polarized sentiment (mostly Negative). The annotators might have treated the Assertion class as the act for Objective sentences, while treating Expression as the act for Subjective sentences. This is arguably easier than sentiment analysis which might explain why annotators agree more on the Speech Act label than the Sentiment label for the ArSAS dataset. Further analysis is required to explain the trends of this dataset.